# FairPO: Robust Preference Optimization for Fair Multi-Label Learning

### v2

@@ 1 Introduction
Multi-label classification (MLC), the task of assigning a relevant subset of labels from a universe $\mathcal{T}$ to an instance $x$, is pervasive in modern machine learning, from image annotation (Wei et al., 2016) to document categorization (Kiritchenko and Matwin, 2006) and bioinformatics (Schietgat et al., 2010). Standard MLC models, parameterized by $\mathbf{w}$, are often trained by minimizing an aggregate loss, such as the sum of per-label binary cross-entropies (BCE) over a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ (Zhang and Zhou, 2013; Sorower, 2010; Tsoumakas and Katakis, 2007). However, such global optimization can obscure or even exacerbate significant fairness issues (Mehrabi et al., 2021; Chouldechova, 2017). Real-world label sets frequently include categories with imbalanced frequencies, differing societal importance, or varying inherent difficulty levels, which can result in models with disparate performance across labels (Zhang and Wu, 2021). For instance, a system might excel on common labels while consistently failing on rare but critical ones. This disparity is particularly problematic when certain labels correspond to protected attributes or groups, demanding equitable treatment.
@@
A key challenge lies in the model’s ability to discriminate effectively, especially for "hard" examples. Standard losses like BCE (Ruby and Yendapalli, 2020) may provide insufficient signal for nuanced distinctions. For a given label $l$, if it is truly positive ($y_{il}=+1$), the model might still assign a low score $m(x_{i};\mathbf{w}_{l})$ or, more critically, assign a higher score to an incorrect label $k$ such that $y_{ik}=0$, i.e., a ‘confusing negative’ that the model mistakenly ranks above the true label. Conversely, if label $l$ is truly negative ($y_{il}=0$), the model might erroneously assign it a high score $m(x_{i};\mathbf{w}_{l})$ (a confusing positive). Existing fairness interventions often target single-label settings (Zafar et al., 2017; Hardt et al., 2016; Dwork et al., 2012) and are not readily adapted to the multi-label context and these symmetric discriminative challenges.
@@
To address this, we introduce FairPO (Fair Preference Optimization), a novel framework that integrates preference-based learning with group robust optimization. Inspired by the success of Direct Preference Optimization (DPO) in aligning models with human preferences (Rafailov et al., 2023), FairPO recasts parts of the MLC task as learning explicit preferences. We partition the label set $\mathcal{T}$ into a privileged set $\mathcal{P}$, where enhanced, fair performance is paramount, and a non-privileged set $\bar{\mathcal{P}}=\mathcal{T}\setminus\mathcal{P}$, where the goal is to maintain robust baseline performance.
Figure 1: FairPO’s methodology: Labels are split into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$). Preference loss ($\mathcal{L}_{\mathcal{P}}$) targets $\mathcal{P}$ labels; constrained loss ($\mathcal{L}_{\bar{\mathcal{P}}}$) maintains $\bar{\mathcal{P}}$ performance. GRPO adaptively balances these for fairer, robust MLC.
For all the privileged labels $l\in\mathcal{P}$, FairPO employs a conditional objective. If a confusing set exists for $(x_{i},l)$—meaning there are confusing negatives $k\in S_{il}^{\text{neg}}$ (where $y_{ik}=0,m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=+1$) or confusing positives $k^{\prime}\in S_{il}^{\text{pos}}$ (where $y_{ik^{\prime}}=+1,m(x_{i};\mathbf{w}_{k^{\prime}})\leq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=0$)—a preference loss (inspired by DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), or CPO (Xu et al., 2024)) is applied. This loss encourages $m(x_{i};\mathbf{w}_{l})\gg m(x_{i};\mathbf{w}_{k})$ for the $(l,k)$ pair if $y_{il}=+1$, or $m(x_{i};\mathbf{w}_{l})\ll m(x_{i};\mathbf{w}_{k^{\prime}})$ for the $(l,k^{\prime})$ pair if $y_{il}=0$. If no such confusing examples are found for $(x_{i},l)$, a standard BCE loss for label $l$ is applied to ensure consistent learning. For non-privileged labels $j\in\bar{\mathcal{P}}$, FairPO enforces a constraint that their classification loss $\ell(\mathbf{w}_{j};x_{i},y_{ij})$ does not substantially exceed that of a reference model $\hat{\mathbf{w}}_{j}$, i.e., $\ell(\mathbf{w}_{j})\leq\ell(\hat{\mathbf{w}}_{j})+\epsilon$.
@@
To manage the trade off between these two distinct objectives, FairPO leverages a group specific robust optimization technique—Group Robust Preference Optimization (GRPO) (Ramesh et al., 2024). The overall learning problem is formulated as a standard minimax optimization: $\min_{\{\mathbf{w}_{t}\}_{t\in\mathcal{T}}}\max_{\alpha\in\Delta_{1}}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}\right]$, where $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$ are the aggregate losses for the privileged and non-privileged groups respectively, and $\alpha=(\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}})$ are adaptive weights in the simplex $\Delta_{1}$. This dynamically balances training, preventing performance degradation in one group for gains in another, thereby promoting robust fairness across the label partition (see Figure 1).
@@
Our contributions are: (1) FairPO, a novel framework for fair MLC using preference signals and group robustness; (2) A conditional, preference-based objective for privileged labels targeting both confusing negatives and confusing positives, with a fallback to standard classification signals; (3) A robust optimization strategy to balance performance across privileged and non-privileged groups. Our comprehensive experiments, evaluating DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), and CPO (Xu et al., 2024) inspired preference objectives within FairPO, demonstrate its efficacy. Notably, FairPO, particularly the CPO-variant, achieves significant improvements in mAP and Exact Match Ratio for privileged labels—for instance, gaining up to 3.44% mAP on COCO (Lin et al., 2015) for the least frequent 20% of labels—while robustly preserving the performance metrics on non-privileged labels over the standard baseline. This underscores FairPO’s ability to effectively address challenging fairness-performance trade-offs in MLC. In Appendix A, we also briefly touch upon the framework’s applicability to fair multi-label attribute generation.
2 Related Work
Fairness in Multi-Label Classification: Ensuring fairness in MLC (Zhang and Wu, 2021; Tantithamthavorn et al., 2018) is complex due to multi-faceted label relationships. Recent efforts address MLC-specific fairness challenges, such as tackling label imbalance impacting tail labels (Wu et al., 2023), learning instance and class-level subjective fairness (Gao et al., 2023), or incorporating fairness in dynamic learning settings like class-incremental MLC (Masud et al., 2023). FairPO contributes by explicitly partitioning labels into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets, applying distinct fairness-motivated objectives to each—notably using preference signals for $\mathcal{P}$—and managing them via a robustness framework. This targeted approach to enhancing performance for pre-defined critical labels, while safeguarding others, differentiates our work.
Preference Optimization: Preference optimization, especially for aligning LLMs (Ouyang et al., 2022; Christiano et al., 2017), has rapidly advanced. Direct Preference Optimization (DPO) (Rafailov et al., 2023) and its reference-free variants like CPO (Xu et al., 2024) and SimPO (Meng et al., 2024) optimize policies directly from preference pairs. The field continues to evolve with methods like Identity Preference Optimization (IPO) for stability (Azhar et al., 2023), Kahneman-Tversky-based optimization (KTO) (Ethayarajh et al., 2024), simple yet effective Rejection Sampling Fine-Tuning (RFT) (Dong et al., 2023), and ongoing theoretical analyses of these methods (Zhao et al., 2024). We use preferences not to rank entire outputs, but to specifically differentiate true label scores from those of their confusing positive/negative counterparts within the privileged label set, thereby sharpening the model’s decision boundaries for critical distinctions.
Group Robust Optimization: Distributionally Robust Optimization (DRO) (Ben-Tal et al., 2009), particularly Group DRO (Sagawa et al., 2019), aims to improve worst-case performance across predefined data groups, enhancing fairness and robustness. This concept sees continued development, for instance, in improving its practicality (Zhang et al., 2022). Group Robust Preference Optimization (GRPO) (Ramesh et al., 2024) extended this to LLM preference learning, balancing performance across preference groups. FairPO directly employs GRPO’s adaptive optimization strategy. However, our groups are defined by the label partition ($\mathcal{P}$ and $\bar{\mathcal{P}}$), and GRPO balances their distinct, custom-formulated loss objectives. This provides a principled mechanism for robustly managing the specific fairness-performance trade-offs in our MLC context.
@@ 3 Preliminaries: Preference Optimization Methods
Direct Preference Optimization (DPO): DPO (Rafailov et al., 2023) directly optimizes a policy $\pi_{\theta}$ using preference pairs $(x,y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$ for prompt $x$. Assuming a Bradley-Terry preference model tied to an implicit reward function related to $\pi_{\theta}$ and a reference policy $\pi_{\text{ref}}$, DPO maximizes the likelihood of observed preferences, resulting in the loss:
@@ 4.1 Problem Setup and Fairness Goals
Given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ is an input instance and $y_{i}=[y_{i1},\dots,y_{iT}]$ is the ground truth binary vector for $T$ labels, we aim to learn a model that assigns a score $m(x_{i};\mathbf{w}_{t})$ for each label $t\in\mathcal{T}$, using parameters $\mathbf{w}_{t}$. These parameters are fine-tuned from a pretrained policy $\pi_{\theta}$, and we leverage a reference model with parameters $\hat{\mathbf{w}}_{t}$ producing scores $m(x_{i};\hat{\mathbf{w}}_{t})$. In practice, a pretrained model $\pi_{\theta}$ (e.g., an LLM) yields an input representation $\mathbf{z}_{i}=\pi_{\theta}(x_{i})\in\mathbb{R}^{d}$. A classifier head (e.g., FFN) with parameters $\mathbf{w}_{t}$ then computes the score $m(x_{i};\mathbf{w}_{t})$ for label $t$. The reference parameters $\hat{\mathbf{w}}$ are typically obtained from standard supervised fine-tuning (SFT) of $\pi_{\theta}$ on $\mathcal{D}$.
Our fairness goals are as follows: For any label $l\in\mathcal{P}$ (Privileged Group), the model should accurately discriminate its true state ($y_{il}=+1$ or $y_{il}=0$) from confusing alternatives. For any label $j\in\bar{\mathcal{P}}$ (Non-Privileged Group), the classification performance under $\mathbf{w}_{j}$ should remain close to that of $\hat{\mathbf{w}}_{j}$. To formalize the challenges for privileged labels, we define two types of confusing set for an instance $x_{i}$ and a privileged label $l\in\mathcal{P}$: If $y_{il}=+1$ (true positive), $S_{il}^{\text{neg}}=\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})\}$ is the set of confusing negatives. If $y_{il}=0$ (true negative), $S_{il}^{\text{pos}}=\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k})\leq m(x_{i};\mathbf{w}_{l})\}$ is the set of confusing positives. The overall confusing set for $(x_{i},l)$ is $S_{il}=S_{il}^{\text{neg}}\cup S_{il}^{\text{pos}}$.
@@ 4.2 Objective for Privileged Labels ($l\in\mathcal{P}$)
Confusing Examples Exist ($S_{il}\neq\emptyset$): We employ a DPO-inspired (Rafailov et al., 2023) preference loss. If $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$ is sampled, we prefer $l$ over $k$ ($l\succ k$). The DPO term is:
@@
The loss contribution is $-\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},l,k))$.
@@
The loss contribution is $-\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},k,l))$.
@@
$\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\},\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\})$ $=\mathbb{E}_{(x_{i},l)\text{ s.t. }l\in\mathcal{P}}\left[\mathbf{1}_{S_{il}\neq\emptyset}\cdot(-\log\sigma(\beta\cdot h_{\text{DPO}}))\right.$
@@
where $(m(x_{i};\text{preferred}),m(x_{i};\text{dispreferred}))$ are $(m(x_{i};\mathbf{w}_{l}),m(x_{i};\mathbf{w}_{k}))$ if $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$, or $(m(x_{i};\mathbf{w}_{k}),m(x_{i};\mathbf{w}_{l}))$ if $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$. The term $\beta\log(\cdot)$ captures the relative preference, while $-\gamma$ enforces a desired separation margin. This $L_{\text{pref}}^{\text{SimPO}}$ replaces the DPO term in Eq. 11 when $S_{il}\neq\emptyset$, with the BCE fallback (Eq. 10) for $S_{il}=\emptyset$ remaining. More broadly, this adaptation leverages SimPO’s margin mechanism for more distinct score separation in challenging MLC cases.
@@
$\mathcal{L}_{\mathcal{P}}^{\text{CPO}}(\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\})=\mathbb{E}_{(x_{i},l)\text{ s.t. }l\in\mathcal{P}}\Big{[}$ $\mathbf{1}_{S_{il}\neq\emptyset}\cdot L_{\text{pref}}^{\text{CPO}}(x_{i},l,k_{s})+\lambda_{\text{CPO}}\cdot\ell_{\text{BCE}}(x_{i},y_{il};\mathbf{w}_{l})\Big{]},$ (14)

### v3

@@ 1 Introduction
Multi-label classification (MLC)—assigning a subset of labels from a universe $\mathcal{T}$ to an instance xx—is a pervasive task in fields like image annotation (Wang et al., 2016) and document categorization (Zangari et al., 2024; Schietgat et al., 2010). Typically, models are trained by minimizing per-label binary cross-entropies (BCE) over a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ (Zhang and Zhou, 2014; Sorower, 2010; Tarekegn et al., 2024). However, this global optimization often creates fairness issues (Mehrabi et al., 2022a; Chouldechova, 2017), as imbalanced label frequencies and varying importance lead to disparate performance (Mehrabi et al., 2022b). For instance, a model might excel on common labels while consistently failing on rare but critical ones. This disparity is particularly problematic when certain labels correspond to protected attributes or groups, demanding equitable treatment.
Figure 1: FairPO’s methodology
@@
A key challenge lies in the model’s ability to discriminate effectively, especially for hard examples. Standard losses like BCE (Ruby and Yendapalli, 2020) may provide insufficient signal for nuanced distinctions. For a given label $l$, if it is truly positive ($y_{il}=+1$), the model might still assign a low score $m(x_{i};\mathbf{w}_{l})$ or, more critically, assign a higher score to an incorrect label $k$ such that $y_{ik}=0$, i.e., a confusing negative that the model mistakenly ranks above the true label. Conversely, if label $l$ is truly negative ($y_{il}=0$), the model might erroneously assign it a high score $m(x_{i};\mathbf{w}_{l})$ (a confusing positive). Existing fairness interventions often target single-label settings (Zafar et al., 2017; Hardt et al., 2016; Dwork et al., 2011) and are not readily adapted to the multi-label context and these symmetric discriminative challenges.
@@
To address this, we introduce FairPO (Fair Preference Optimization), a novel framework that integrates preference-based learning with group robust optimization. Inspired by the success of Direct Preference Optimization (DPO) in aligning models with human preferences (Rafailov et al., 2024), FairPO recasts parts of the MLC task as learning explicit preferences. We partition the label set $\mathcal{T}$ into a privileged set $\mathcal{P}$, where enhanced, fair performance is paramount, and a non-privileged set $\bar{\mathcal{P}}=\mathcal{T}\setminus\mathcal{P}$, where the goal is to maintain robust baseline performance.
@@
For all the privileged labels $l\in\mathcal{P}$, FairPO employs a conditional objective. If a confusing set exists for $(x_{i},l)$—meaning there are confusing negatives $k\in S_{il}^{\text{neg}}$ (where $y_{ik}=0,m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=+1$) or confusing positives $k^{\prime}\in S_{il}^{\text{pos}}$ (where $y_{ik^{\prime}}=+1,m(x_{i};\mathbf{w}_{k^{\prime}})\leq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=0$)—a preference loss (inspired by DPO (Rafailov et al., 2024), SimPO (Meng et al., 2024), or CPO (Xu et al., 2024)) is applied. This loss encourages $m(x_{i};\mathbf{w}_{l})\gg m(x_{i};\mathbf{w}_{k})$ for the $(l,k)$ pair if $y_{il}=+1$, or $m(x_{i};\mathbf{w}_{l})\ll m(x_{i};\mathbf{w}_{k^{\prime}})$ for the $(l,k^{\prime})$ pair if $y_{il}=0$. If no such confusing examples are found for $(x_{i},l)$, a standard BCE loss for label $l$ is applied to ensure consistent learning. For non-privileged labels $j\in\bar{\mathcal{P}}$, FairPO enforces a constraint that their classification loss $\ell(\mathbf{w}_{j};x_{i},y_{ij})$ does not substantially exceed that of a reference model $\hat{\mathbf{w}}_{j}$, i.e., $\ell(\mathbf{w}_{j})\leq\ell(\hat{\mathbf{w}}_{j})+\epsilon$. To manage the trade off between these two distinct objectives, FairPO leverages a group specific robust optimization technique—Group Robust Preference Optimization (GRPO) (Ramesh et al., 2024). The overall learning problem is formulated as a standard minimax optimization: $\min_{\{\mathbf{w}_{t}\}_{t\in\mathcal{T}}}\max_{\alpha\in\Delta_{1}}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}\right]$, where $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$ are the aggregate losses for the privileged and non-privileged groups respectively, and $\alpha=(\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}})$ are adaptive weights in the simplex $\Delta_{1}$. This dynamically balances training, preventing performance degradation in one group for gains in another, thereby promoting fairness across the labels (see Figure 1).
@@
Our primary contributions are a novel framework for fair MLC using preference signals, a conditional objective targeting hard examples, and a robust optimization strategy to manage fairness trade-offs. Our comprehensive experiments show that FairPO, particularly the CPO-variant, achieves significant gains—for instance, up to 3.44% mAP on the least frequent labels of COCO (Lin et al., 2015)—while robustly maintaining performance on non-privileged labels. We also outline the framework’s applicability to attribute generation in Appendix A.
@@ 2 Preliminaries: Preference Optimization Methods
Direct Preference Optimization (DPO): DPO (Rafailov et al., 2024) directly optimizes a policy $\pi_{\theta}$ using preference pairs $(x,y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$ for prompt $x$. Assuming a Bradley-Terry preference model tied to an implicit reward function related to $\pi_{\theta}$ and a reference policy $\pi_{\text{ref}}$, DPO maximizes the likelihood of observed preferences, resulting in the loss:
@@ 3.1 Problem Setup and Fairness Goals
Given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, we fine-tune parameters $\mathbf{w}_{t}$ for a per-label classifier that assigns a score $m(x_{i};\mathbf{w}_{t})$ for each label $t\in\mathcal{T}$. Our framework leverages a reference model with parameters $\hat{\mathbf{w}}_{t}$, typically obtained from standard supervised fine-tuning (SFT).
Our framework addresses the problem of group fairness, where the goal is to ensure equitable performance across different predefined groups. In our MLC context, we define these groups not by instance attributes (e.g., demographics), but by partitioning the labels themselves into a privileged set $\mathcal{P}$ (e.g., rare, critical, or historically underperforming labels) and a non-privileged set $\bar{\mathcal{P}}$. Our fairness objective is a nuanced form of equitable treatment: we aim to significantly improve performance for the underperforming privileged group while ensuring that performance for the non-privileged group is not substantially harmed. This goal of targeted improvement without undue harm to other groups is enforced by our robust optimization framework.
For any privileged label $l\in\mathcal{P}$, the model must accurately discriminate its true state from confusing alternatives. For any non-privileged label $j\in\bar{\mathcal{P}}$, performance should not degrade significantly from the reference model. To formalize this, we define two types of confusing sets for a privileged label $l\in\mathcal{P}$: the set of confusing negatives, $S_{il}^{\text{neg}}=\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})\}$ when $y_{il}=+1$, and the set of confusing positives, $S_{il}^{\text{pos}}=\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k})\leq m(x_{i};\mathbf{w}_{l})\}$ when $y_{il}=0$. The overall confusing set is $S_{il}=S_{il}^{\text{neg}}\cup S_{il}^{\text{pos}}$.
@@ 3.2 Objective for Privileged Labels ($l\in\mathcal{P}$)
Confusing Examples Exist ($S_{il}\neq\emptyset$): We employ a DPO-inspired (Rafailov et al., 2024) preference loss. If $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$ is sampled, we prefer $l$ over $k$ ($l\succ k$). The DPO term is:
@@
The loss contribution is $\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},l,k))$.
@@
The loss contribution is $\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},k,l))$.
@@
$\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t}|t\in\mathcal{T}\},\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\})$ $=\mathbb{E}_{(x_{i},l)\text{ s.t. }l\in\mathcal{P}}\left[\mathbf{1}_{S_{il}\neq\emptyset}\cdot(-\log\sigma(\beta\cdot h_{\text{DPO}}))\right$.
@@
where $(m(x_{i};\text{preferred}),m(x_{i};\text{dispreferred}))$ are $(m(x_{i};\mathbf{w}_{l}),m(x_{i};\mathbf{w}_{k}))$ if $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$, or $(m(x_{i};\mathbf{w}_{k}),m(x_{i};\mathbf{w}_{l}))$ if $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$. The term $\beta\log(\cdot)$ captures the relative preference, while $\gamma$ enforces a desired separation margin. This $L_{\text{pref}}^{\text{SimPO}}$ replaces the DPO term in Eq. 11 when $S_{il}\neq\emptyset$, with the BCE fallback (Eq. 10) for $S_{il}=\emptyset$ remaining. More broadly, this adaptation leverages SimPO’s margin mechanism for more distinct score separation in challenging MLC cases.
@@
[…]